OpenStudy (anonymous):

Find the derivatives of the following question w.r.t.x tan^-1(secx+tanx)

OpenStudy (gorv):

u know what is derivative of tan^(-1)x???

OpenStudy (gorv):

\[\huge D(\tan^{-1} x)=\huge \huge \frac{ 1 }{ 1+x^2 }\]

OpenStudy (anonymous):

Yes I know..

OpenStudy (gorv):

so here x= secx+tanx

OpenStudy (gorv):

\[=\huge \frac{ 1 }{ 1+(scex+tanx) ^2} * D(secx+tanx)\]

OpenStudy (gorv):

can u do it now ??

OpenStudy (gorv):

what is derivative of secx ??

OpenStudy (anonymous):

Secx*tanx

OpenStudy (gorv):

and that of tanx ??

OpenStudy (anonymous):

Sec^2x

OpenStudy (gorv):

sou did it thatz itt

OpenStudy (gorv):

\[\large \frac{ secx*tanx+\sec^2x }{ 1+(secx+tanx)^2 }\]

OpenStudy (anonymous):

is it the total answer..??

OpenStudy (gorv):

yeah ....

OpenStudy (anonymous):

but the answer is 1/2

OpenStudy (gorv):

u can do it further ..by expanding the square

OpenStudy (gorv):

numerator

OpenStudy (gorv):

keep it as it is ..and come to denominator first

OpenStudy (gorv):

1+sec^2+tan^2+2*tan*sec

OpenStudy (gorv):

when we open the square bracket

OpenStudy (anonymous):

sir can you draw it in equation..i am bit confused

OpenStudy (gorv):

lol i m no sir

OpenStudy (anonymous):

you are teaching me so ofcourse you are a sir to me

OpenStudy (gorv):

\[1+(secx+tanx)^2=1+(\sec^2x+\tan^2x+2tanx*secx)\]

OpenStudy (gorv):

u got it now ..we are solving denominator first

OpenStudy (anonymous):

check my solution

OpenStudy (anonymous):

@lanku

OpenStudy (anonymous):

@Princer_Jones My question is a bit different..!!

OpenStudy (gorv):

we need to find derivative

OpenStudy (anonymous):

Refer to the attachment for some surprising results.

OpenStudy (anonymous):

Yes just find the derivative it is 1/2 i gave the expression to it.

OpenStudy (anonymous):

@gorv after solving denominator how to proceed to algebra

OpenStudy (anonymous):

arc tan means tan inverse

OpenStudy (anonymous):

@lanku i already did the solution

OpenStudy (anonymous):

tan inverse tan theta =theta.

OpenStudy (anonymous):

@Princer_Jones main problem is there that I don't know how I will simplify it..

OpenStudy (gorv):

\[1+\sec^2x+\tan^2x+2tanx*\]

OpenStudy (gorv):

we know that \[\sec^2x-\tan^2x=1\]

OpenStudy (gorv):

replace one by this ??

OpenStudy (anonymous):

sir can you do it from first to last step by step

OpenStudy (gorv):

\[\sec^2x-\tan^2x+\sec^2x+\tan^2x+2tanx*secx\]

OpenStudy (anonymous):

@gorv @lanku now check the attachment

OpenStudy (gorv):

lol dont call me sir

OpenStudy (gorv):

@lanku u can go by his approch also that is easy

OpenStudy (anonymous):

ok ok but I still don't get it..please do it from first to last if you have time

OpenStudy (anonymous):

I did everything. you just tell me which step you dont understand

OpenStudy (anonymous):

after that i don't get anything

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

I did in a different way, You check my attachment,

OpenStudy (anonymous):

ok..thank you

OpenStudy (anonymous):

check the attachment, it is the shortest method

OpenStudy (anonymous):

but sorry to say I didn't understand that.... where the tan^1 (Secx+tanx) gone??

OpenStudy (anonymous):

we calculated sec x+tan x as tan(x/4+pi/2) and you know that tan inverse tan theta is theta.

OpenStudy (anonymous):

so the given expression reduces to x/2+pi/4, whose derivative is 1/2

OpenStudy (anonymous):

@lanku

OpenStudy (anonymous):

yeah i am getting a bit...let me try once..!

OpenStudy (anonymous):

ok,,

OpenStudy (anonymous):

do i have to assume x=? for this equation to get theta ?

OpenStudy (anonymous):

no need of assuming, you just write the steps i wrote in the attachment, first compute tan x+sec x=tan (x/2+pi/4) as i proved it. then just tan inverse (sec x+tan x)=x/2+pi/4

OpenStudy (anonymous):

actually our study module is different than yours thatswhy I am getting too much of confusion...but its owk i will find it..thanks for your answer..!!

OpenStudy (anonymous):

@lanku I always try to do calculus in methods which are not given in the book. But many times the calculation of 5 minutes is reduced to calculation of 30seconds.Hence I choose always a different way